OpenStudy (anonymous):

What is the solution of the equation? 16 = –d + 6 A. 10 B. –10 C. –9 D. –15 need help i want to show my son how to do it but i failed in math so can someone run me through this so i can show him thanks

OpenStudy (anonymous):

remember when moving a term from one side of the equation to the other ,you change its sign 16=-d+6 16+d=6 d=6-16 d=-10

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

he had 10 but not negative

OpenStudy (anonymous):

no if 10 then you;ll get 16=-10+6=>16=-4 wrong if -10 16=-(-10)+6=>16=10+6=>16=16 right..